The second hole of the site was spudded with the APC at 0650 hr on 28 August. The seafloor depth inferred from recovery was 544.2 mbsl. APC coring advanced to 68.3 mbsf when Core 181-1121B-8H did not achieve a full stroke. Cores were oriented starting with 181-1120B-3H. Coring was continued with the XCB and deepened the hole to 188.0 mbsf when operations were put on standby because of excessive heave (2 m with an occasional 3 m event) generated by 2.02.5-m seas combined with a 4.05.5 m swell from the southwest. The hole was abandoned, the bit clearing the seafloor at 2020 hr on 28 August.
A force 9 gale with sustained winds up to 45 kt and gusts to 60 kt (force 10) prevailed in the drilling area from the evening of 28 August until 0045 hr on 30 August when it was decided to continue site operations even through environmental conditions were borderline for shallow-water operations. After positioning the bit at 543 mbsl, Hole 1120C was spudded with the APC at 0613 hr on 30 August. The seafloor depth based upon recovery was 545.9 mbsl. Piston coring advanced to 44.6 mbsf when an incomplete stroke was encountered at Core 181-1120C-5H. While recovering the core, the wireline broke. In an attempt to fish the parted wireline, monel sinker bars, and APC assembly, the drill pipe was pulled back. The bit cleared the seafloor at 0613 hr on 31 August.
A wireline spear was made up and run in the pipe. It tagged the broken wireline at 183 mbsl. Attempts to work the broken wireline back up the pipe were unsuccessful and the spear slipped off. The drill pipe was pulled back in stands until the broken end of the wireline was found. After the broken wireline was recovered, the APC core barrel containing Core 181-1120C-5H was retrieved at 1055 hr. Environmental conditions were too extreme to continue and operations were placed in standby mode waiting on weather to improve.
At 1600 hr on 31 August, Hole 1120D was spudded and drilled ahead to 157.4 mbsf where XCB coring was initiated. XCB coring advanced from 157.4 to 220.7 mbsf when operations again had to be terminated because of excessive heave. It was decided to discontinue coring at this location and move on to Site 1121. As the drill string was being retrieved, the beacon was recalled and recovered. At 0600 hr on 1 September, the vessel began the transit to Site 1121.
To 181 Summary of Engineering and Drilling Operations: Site 1121
To 181 Table of Contents